OpenStudy (anonymous):

Let f(x) =( 1/(x^3-2x)) . Then f' =

OpenStudy (yttrium):

You know quotient rule?

OpenStudy (anonymous):

i havent even learned that yet

OpenStudy (yttrium):

Don't you know the formula (vdu-udv)/v^2 ???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

nominate a "u" and a "v" u = 1 v = x^3-2x

OpenStudy (anonymous):

then derive both: u' = 0 v' = x^3 - 2x where you have to use product rule... for v' do you know that rule?

OpenStudy (anonymous):

u'v + v'u

OpenStudy (anonymous):

^^ product rule

OpenStudy (anonymous):

meh, i'll try it

OpenStudy (anonymous):

u = x^3 v = -2x so: u' = 3x^2 v' = -2

OpenStudy (anonymous):

so using product rule: u'v + v'u 3x^2(-2x) -2(x^3)

OpenStudy (anonymous):

simplified down: -6x^3-2x^3

OpenStudy (anonymous):

so using quotient rule: (vu'-uv')/v^2